# Software Framework for Optically Accessible Quantum Memory Using Group-IV Color Centers in Diamond

### v2

@@ III.1 Single-Photon Source
with the normalization factor $\mathcal{N}=\left(\int_{\mathbb{R}}|S(\omega)|^{2}\right)^{-1/2}$ and the bandwidth $\gamma$.
An arbitrary photonic qubit is then given by
$|\psi_{\rm ph}\rangle=\alpha|e\rangle+\beta|l\rangle,\quad|\alpha|^{2}+|\beta|^{2}=1.$ (8)
We assume that the emitted photon gets depolarized during its generation according to the quantum channel [53]
@@
with $\epsilon=2(1-F)$ and $\rho_{\rm ph}=|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|$ where the quantity $F$ denotes the mixed state fidelity between the states $\rho_{\rm ph}$ and $\mathcal{E}(\rho_{\rm ph})$ [54] .
@@ III.2 Cavity Design
To store a photon with a finite bandwidth in the cavity coupled spin we must carefully engineer the cavity such that the controlled phase gate fidelity between the spin and photon [54]
$F_{\rm sp}=\langle\rm Bell|\rho_{\rm sp}|\rm Bell\rangle$ (10)
with $|\rm Bell\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized (see App. B). To evaluate the fidelity given in Eq. (82), the spin-dependent reflection coefficients are required. We assume that a single cavity mode is coupled to the $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ transitions. The reflection coefficients are given by [55, 49]
@@
with the cavity loss rate $\kappa$, the detunings $\Delta_{1A}:=\omega-\omega_{1A},\Delta_{2B}=\omega-\omega_{2B}$, the transition frequencies $\omega_{1A},\omega_{2B}$, the atomic decay rates $\gamma_{1A},\gamma_{2B}$, the detuning $\Delta_{c}=\omega-\omega_{c}$ with the cavity mode central frequency of the SnV $\omega_{c}$ and the coupling strength between the SnV and the photonic mode in the cavity
@@
where the mode volume is $V=V_{\rm eff}\frac{\lambda^{3}}{2n^{3}}$ [56], the wavelength of the cavity mode $\lambda=2\pi c/\omega_{c}$, the speed of light $c$ and the mode orientation $\bm{\epsilon}$. We assume the effective mode volume $V_{\rm eff}=1.8$ and mode orientation $\bm{\epsilon}=\bm{e_{z}}$, i.e. the mode is parallel to the SnV’s symmetry axis [6].
To maximize the fidelity shown in Eq. (82) the software package can optimize the triple $(\omega_{0},\omega_{c},\kappa)$ using simplicial homology global optimization [57, 58]. Optimized values for $(\omega_{c},\kappa)$ yield a cooperativities $\mathbf{C}=(C_{1A},C_{2A},C_{1B},C_{2B})$ calculated by $C_{kl}=2|g_{kl}|^{2}/\kappa\gamma_{kl}$ for $k=1,2$, $l=A,B$, which is necessary for simulating the influence of the Purcell enhancement [59].
@@ III.3 Spin Rotation
with the free contribution $H_{0}$ and the light-matter interaction term $\bm{\mu}\cdot\mathbf{E}(t)$ using the optimized parameters for $B_{\rm dc}=3.0$ T and $\theta_{\rm dc}=43.11$ deg as demonstrated in [6]. The detailed description of the Hamiltonian $H^{\rm opt}(t)$ and temperature, cooperativity and magnetic field dependent relaxation rates $\gamma_{k}$ as well as Lindblad-operators $\{L_{k}\}$ modeling photonic decay processes using Fermi’s golden rule and fast phononic processes are explained in [47, 6].
@@
consisting of the time-invariant part $H_{\rm dc}$ due to the DC magnetic field and the driving term $H_{\rm ac}(t)$ modeling an oscillating AC magnetic field and the dissipator $\mathcal{L}_{D}^{(T,E_{x},\epsilon_{xy},\bm{B_{\rm dc}})}(\rho)$ containing only the temperature, strain and field $\bm{B_{\rm dc}}$ dependent phononic processes. The $\pi/2$ rotation is achieved by driving the AC field for a quarter of a full Rabi oscillation [45].
@@ III.4 Kraus-Operators
To evaluate Kraus-operators for the read-in and read-out process we need to propagate the full basis of photonic - and spin states subject to the quantum channel for the read-in and read-out process, respectively. A detailed mathematical exploration of the spin states after measurement including imperfect spin gates using microwave or optical control encountering photonic and fast phononic processes and a detailed spin-photon interaction model including crosstalk which goes beyond the work in [49] is outlined in the Apps. A, C. Let’s assume a photonic qubit $\rho_{\rm ph}$ is passing all reflection steps including the $X$-measurement resulting in the measurement outcomes $\rho_{+}$ and $\rho_{-}$.
@@
$\mathcal{D}_{\rm read-in}(|i\rangle\langle j|)=\rho_{\rm sp}^{ij}(T_{\rm read-in}),\quad i,j=e,l$ (25)
with the quantum channel for the read-in $\mathcal{D}_{\rm read-in}$ assuming a measurement in $|+\rangle$ for the photonic qubit.
To compute Kraus-operators for the read-out process we evaluate
$\mathcal{D}_{\rm read-out}(|i\rangle\langle j|)=\rho_{\rm ph}^{ij}(T_{\rm read-out}),\quad i,j=1,2$ (26)
with the quantum channel for the read-out $\mathcal{D}_{\rm read-out}$ assuming a measurement in the state $|1\rangle$ for the spin qubit. Details regarding the read-out process are given in the App. D. We compute Kraus-operators by evaluating Eq. (24). For a photonic qubit of the form shown in Eq. (9) the spin state after read-in is
$\rho_{\rm sp}=\frac{\mathcal{D}_{\rm read-in}(\rho_{\rm ph})}{{\rm tr}(\mathcal{D}_{\rm read-in}(\rho_{\rm ph}))}$ (27)
and after read-out we get
$\rho_{\rm ph}=\frac{\mathcal{D}_{\rm read-out}(\rho_{\rm sp})}{{\rm tr}(\mathcal{D}_{\rm read-out}(\rho_{\rm sp}))}.$ (28)
@@ III.5 Approximation Error
where $\left\lVert A\right\rVert_{1}=\max_{1\leq j\leq 8}\left(\sum_{i=1}^{8}|A_{ij}|\right)$ denotes the 1-norm of a matrix. We also use Eq. (29) for evaluating the approximation error for microwave control.
@@ IV Numerical Examples
[…]

### v3

@@ III.1 Single-Photon Source
with the normalization factor $\mathcal{N}=\left(\int_{\mathbb{R}}|S(\omega)|^{2}\right)^{-1/2}$, a field strength $\epsilon_{0}$ and the bandwidth $\gamma=\gamma_{\rm in},\gamma_{\rm out}$ where $\gamma_{\rm in,out}$ denote the bandwidth of the incoming photon for the read-in and read-out process, respectively. An arbitrary photonic qubit is then given by
$|\psi_{\rm ph}\rangle=\alpha|E\rangle+\beta|L\rangle$ (11)
with $\alpha,\beta\in\mathbb{C}$ and $|\alpha|^{2}+|\beta|^{2}=1$. We assume that the emitted photon gets depolarized during its generation according to the quantum channel [64]
@@
with $\epsilon=2(1-F_{\rm ph})$ and $\rho_{\rm ph}=|\psi_{\rm ph}\rangle\langle\psi_{\rm ph}|$ where the quantity $F_{\rm ph}$ denotes the state fidelity between the states $\rho_{\rm ph}$ and $\mathcal{E}(\rho_{\rm ph})$ [65]. For the read-in process we do not include depolarization during photon generation because it is more intuitive for the user to depolarize a photonic qubit according to Eq. (12) and subsequently apply the Kraus map modeling the read-in process. For read-out the photon generation fidelity is called $F_{\rm aux}$ and it holds $\alpha=\beta=1/\sqrt{2}$.
@@ III.2 Cavity Design
To store a broadband photon in the cavity coupled spin we must carefully engineer the cavity such that the controlled phase gate fidelity between the spin and photon [65]
$F_{\rm sp}=\langle\psi|\rho_{\rm sp}|\psi\rangle$ (13)
with $|\psi\rangle=\frac{1}{\sqrt{2}}(|1\rangle+|2\rangle)$ is maximized (see Appendix, Sec. 2 for details). To evaluate the fidelity given in Eq. (13), the spin-dependent reflection coefficients are required. We assume that a single cavity mode is coupled to the $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ transitions. The reflection coefficients are given by [58, 56]
@@
with the detunings $\Delta_{1A}:=\omega-\omega_{1A},\Delta_{2B}=\omega-\omega_{2B}$, the transition frequencies $\omega_{1A},\omega_{2B}$, the atomic decay rates $\gamma_{1A},\gamma_{2B}$, the detuning $\Delta_{c}=\omega-\omega_{c}$ with the cavity mode central frequency of the SiV-/SnV- $\omega_{c}$ and the coupling strength between the SiV-/SnV- and the photonic mode in the cavity
@@
where the mode volume is $V=V_{\rm eff}\frac{\lambda^{3}}{2n^{3}}$ [59], the wavelength of the cavity mode $\lambda=2\pi c/\omega_{c}$, the speed of light $c$ and the mode orientation $\bm{\epsilon}$. We assume the effective mode volume $V_{\rm eff}=1.8$ and mode orientation $\bm{\epsilon}=\bm{e_{z}}$, i.e. the mode is parallel to the SiV’s/SnV’s symmetry axis.
To maximize the fidelity shown in Eq. (13) the software package can optimize the triple $(\omega_{0},\omega_{c},\kappa)$ using simplicial homology global optimization [66, 67]. Optimized values for $(\omega_{c},\kappa)$ yield cooperativities $\mathbf{C}=(C_{1A},C_{2A},C_{1B},C_{2B})$, calculated as $C_{kl}=2|g_{kl}|^{2}/(\kappa\gamma_{kl})$ for $k=1,2$ and $l=A,B$. Care must be taken with the factor of two, since $\kappa$ denotes the half-width at half-maximum (HWHM) of the cavity mode, while $\gamma_{kl}$ refers to the full width at half-maximum (FWHM) of the optical transition. Accordingly, the cooperativity definition consistent with HWHM values should be used when calculating the Purcell enhancement [68].
@@ III.3 Spin Rotation
with the free contribution $H_{0}$ and the light-matter interaction term $\bm{\mu}\cdot\mathbf{E}(t)$ using the optimized parameters for $B_{\rm dc}=0.3,1.0,3.0$ T and $\theta_{\rm dc}$ taken from [6] for the SnV. The detailed description of the Hamiltonian $H^{\rm opt}(t)$ and temperature, cooperativity and magnetic field dependent relaxation rates $\gamma_{k}$ as well as Lindblad-operators $\{L_{k}\}$ modeling photonic decay processes using Fermi’s golden rule and fast phononic processes are explained in [50, 6].
@@
consisting of the time-invariant part $H_{\rm dc}$ due to the DC magnetic field and the driving term $H_{\rm ac}(t)$ modeling an oscillating AC magnetic field and the dissipator $\mathcal{L}_{D}^{(T,E_{x},\epsilon_{xy},\bm{B_{\rm dc}})}(\rho)$ containing only the temperature, strain and field $\bm{B_{\rm dc}}$ dependent phononic processes. The $\pi/2$ rotation is achieved by driving the AC field for a quarter of a full Rabi oscillation for both the SiV- and SnV- [48].
@@ III.4 Kraus-Operators
To evaluate Kraus-operators for the read-in and read-out process we need to propagate the full basis of photonic - and spin states subject to the quantum channel for the read-in and read-out process, respectively. A detailed mathematical exploration of the spin states after measurement including imperfect spin gates using microwave or optical control encountering photonic and fast phononic processes and a detailed spin-photon interaction model including crosstalk which goes beyond the work in [56] is outlined in the Appendix, Sec. 1. Let’s assume a photonic qubit $\rho_{\rm ph}$ is passing all reflection steps including the $X$-measurement resulting in the measurement outcomes $\rho_{+}$ and $\rho_{-}$.
@@
$\mathcal{D}_{\rm read-in,+}(|i\rangle\langle j|)=\rho_{\rm sp,+}^{ij}(T_{\rm read-in}),\quad i,j=E,L,$ (28)
$\mathcal{D}_{\rm read-in,-}(|i\rangle\langle j|)=\rho_{\rm sp,-}^{ij}(T_{\rm read-in}),\quad i,j=E,L$ (29)
with the quantum channel for the read-in $\mathcal{D}_{\rm read-in,+/-}$ assuming a measurement in $|+/$-$\rangle$ for the photonic qubit. To compute Kraus-operators for the read-out process we evaluate
$\mathcal{D}_{\rm read-out,1}(|i\rangle\langle j|)=\rho_{\rm ph,1}^{ij}(T_{\rm read-out}),\quad i,j=1,2,$ (30)
$\mathcal{D}_{\rm read-out,2}(|i\rangle\langle j|)=\rho_{\rm ph,2}^{ij}(T_{\rm read-out}),\quad i,j=1,2$ (31)
with the quantum channel for the read-out $\mathcal{D}_{\rm read-out,1/2}$ assuming a measurement in the state $|1/2\rangle$ for the spin qubit. Details regarding the read-out process are given in the Appendix, Sec. 4. We compute Kraus-operators by evaluating Eq. (27). For a photonic qubit of the form shown in Eq. (12) the spin state after read-in is
$\rho_{\rm sp,+/-}=\frac{\mathcal{D}_{\rm read-in,+/-}(\rho_{\rm ph})}{\eta_{+/-}}$ (32)
with $\eta_{+/-}={\rm tr}(\mathcal{D}_{\rm read-in,+/-}(\rho_{\rm ph}))$ describing the trace reduction of the read-in process depending on the measurement $|+\rangle$ or $|$-$\rangle$. After read-out we get
$\rho_{\rm ph,+/-,1/2}=\frac{\mathcal{D}_{\rm read-out,1/2}(\rho_{\rm sp,+/-})}{\eta_{+/-,1/2}}.$ (33)
with $\eta_{+/-,1/2}={\rm tr}(\mathcal{D}_{\rm read-out,1/2}(\rho_{\rm sp,+/-}))$ representing the trace reduction for the read-out procedure depending on whether the input state was measured in $|+\rangle$ or $|$-$\rangle$ respectively. To evaluate the quantum memory performance using Kraus-operators, we define an average fidelity and trace reduction that account for both the read-in and read-out processes across the Bloch sphere of photonic qubits (see Appendix, Sec. 5 for details).
@@ III.5 Approximation Error
where $\left\lVert A\right\rVert_{1}=\max_{1\leq j\leq 8}\left(\sum_{i=1}^{8}|A_{ij}|\right)$ denotes the 1-norm of a matrix. We also use Eq. (35) for evaluating the approximation error for microwave control. As we will show in Section IV, for errors smaller than $10^{-4}$, the resulting trace reduction can be considered negligible.
@@ IV Numerical Examples
Figure 2: Visualization of the infidelity $1-\langle F\rangle$ and trace $\langle{\rm tr}(\rho_{\rm ph})\rangle$ of the photonic qubit after read-out (see Appendix, Sec. 5 for details) as a function of the system’s nanophotonic temperature $T$ assuming the bandwidth of the photonic qubit for the read-in and read-out process is $\gamma_{\rm in,out}=0.1$ GHz, the AC magnetic field for microwave control of the SiV- (SnV-) is $B_{\rm ac}=3.7(1.0)$ mT, the axial strain is $E_{x}=4\cdot 10^{-5}(0)$ and the shear strain is $\epsilon_{xy}=0(0)$ for the SiV- (SnV-), the fidelity of the auxiliary photon source for read-out is $F_{\rm aux}=1.0$ for all the considered cases and we limit the cooperativity to $C_{\rm max}=12.5$.
Figure 3: a) Visualization of the infidelity as a function of the bandwidth for the SiV- and SnV- for the read-in and b) read-out process. For the SnV- the trace is ${\rm tr}(\rho_{\rm ph})\approx 0.94$ and for the SiV- it is $\langle{\rm tr}(\rho_{\rm ph})\rangle=0.80$ on the considered range. c) Visualization of the infidelity as a function of the maximal allowed cooperativity assuming all microwave control for the SnV- with $B_{\rm dc}=0.3$ T, $B_{\rm ac}=1$ mT in a low strain environment at $T=0.1$ K for $\gamma_{\rm in,out}=1$ GHz. The average trace reduction on that range is $\langle{\rm tr}(\rho_{\rm ph})\rangle\approx 0.9580$. Due to a rather noisy dependence of the trace reduction on the cooperativity we omit that graph here. For the SiV- the optimal cooperativity is below $12.5$ leading to no improvement on the considered range which is why we omit the SiV- representation here. d) Visualization of the infidelity as a function of the infidelity of the auxiliary photon source for read-out considering the SiV- as defect center. The trace is $\langle{\rm tr}(\rho_{\rm ph})\rangle=0.7812$ on the considered range.
The main objective of the software module [49] is the calculation of Kraus operators that describe the action of the memory on a photonic qubit after storage and retrieval. We now provide examples of how to use the module based on the computation of a Kraus representation of the memorization procedure. As an instructive example, we calculate the average fidelity for reading in and reading out a photonic qubit (details are shown in the Appendix, Sec. 5). We also show the trace of the photonic qubit after read-out to quantify loss and decoherence in the system.
We visualize the infidelity and trace as a function of the system’s nanophotonic temperature from $100$ mK to $4$ K, which is the experimentally relevant temperature range for control experiments [52, 53, 54, 55]. To cover the broad range of accessible magnetic field strengths [52, 54], we perform simulations for different fields and for both the SiV and SnV. In addition, we sweep the bandwidth of the single-photon sources to assess how source spectral properties impact overall performance. Since cavity cooperativity is fabrication-limited [60], we also evaluate the influence of the cooperativity bound. Finally, we analyze the memory performance as a function of the fidelity of the auxiliary photon source used for read-out, which is essential for determining the minimum source quality required for practical implementations. As a last example, we show how user-defined cavity parameters can be used to predict the performance of a G4V based quantum memory integrated into a cavity.
In Figure 2 we visualize the infidelity and trace as a function of the temperature for the SiV- and SnV- using three magnetic field strengths and show the results for both microwave and optical control to produce the spin $\pi/2$ rotation. For the simulations we assume a range of experimentally relevant parameters [71, 52, 54]: $\gamma_{\rm in,out}=0.1$ GHz, $B_{\rm ac}=1.0$ mT $(3.7)$ mT for the SiV- (SnV-), strain parameters $E_{x}=4\cdot 10^{-5}$ $(0)$, $\epsilon_{xy}=0$ $(0)$ and the upper bound for the cooperativity $C_{\rm max}=12.5$. The chosen value $C_{\mathrm{max}}=12.5$ does not limit the experimentally achievable fidelity for the SiV- [72], but it does constrain the fidelity for the SnV-, consistent with the cooperativities reported for SnV- cavities in [60]. It becomes apparent that the infidelity increases and the trace decreases as the temperature rises. The infidelity rise is due to phonon induced spin dephasing [48]. Since phononic excitation also plays an increasingly important role as the temperature rises the population leaves the logical subspace leading to a rise of the approximation error and trace drop. To justify the estimated approximation error bound of $10^{-4}$ for negligible trace reduction, we refer to Fig. 2d). In this figure, the approximation error remains below $4\times 10^{-5}$, while the trace shows no noticeable variation on the plotted scale.
In Fig. 2a) and c) we observe that the infidelity is smaller when the DC magnetic field is oriented orthogonally to the symmetry axis than when it is applied parallel to it. This is expected because the orthogonal orientation yields a smaller optical splitting $\Delta\omega_{s}:=\omega_{2B}-\omega_{1A}$ [72]. Comparing the SiV and SnV in these figures, we further see that the SiV exhibits a lower infidelity than the SnV for temperatures below approximately $2$ K, whereas the trend reverses as the temperature approaches $4$ K. The superior low-temperature performance of the SiV arises from the chosen cooperativity limit, which does not reduce the SiV fidelity but does limit the SnV fidelity [72]. At higher temperatures, however, phonon-induced spin dephasing becomes increasingly significant for the spin gate. The microwave spin gate used for the SiV is more sensitive to this effect than the corresponding gate for the SnV. In Fig. 2e) we observe that the infidelity remains essentially independent of temperature when optical control is used for the spin rotation, which follows from the short gate duration relative to the timescale of phonon-induced spin dephasing [6].
In Figure 3 we visualize the infidelity as a function of the bandwidth for read-in and read-out as well as the upper bound on the cooperativity for the SnV and the infidelity of the auxiliary photon source. In Fig. 3a) and b) we plot the infidelity $1-\langle F\rangle$ as a function of the bandwidth $\gamma$ of the incoming and outgoing photonic modes, using either the SiV- or SnV- as the quantum memory. The simulations assume a magnetic field strength $B_{\rm dc}=300$ mT with orientation $\theta_{\rm dc}=\pi/2$, microwave control with $B_{\rm ac}=1$ mT (3.7 mT) and $\theta_{\rm ac}=0$, a temperature of $T=0.1$ K, axial strain $E_{x}=4\cdot 10^{-5}$ (0), shear strain $\epsilon_{xy}=0$ (0), an auxiliary photon-source fidelity of $F_{\rm aux}=1.0$, and a cooperativity bound of $C_{\rm max}=12.5$. As expected, the infidelity increases with bandwidth. This trend reflects the reduced phase contrast between the transitions $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ as the photonic spectral width broadens [56].
When comparing Figure 3a) and b) we observe that the infidelities for read-in and read-out are equal for the SiV-. The same behavior becomes apparent for the SnV-. That is comprehensible because both the read-in and read-out process entail the same scheme for producing entanglement, hence the same errors are produced under an equal parameter setting. We also observe that the SiV’s infidelity is smaller than the SnV’s infidelity on the considered range. That is due to the cooperativity limit $C_{\rm max}=12.5$. While the optimal cooperativity for the SiV- for $\gamma_{\rm in,out}=1$ GHz is below $C_{\rm max}$ it is not for the SnV-. No source of error for the SiV- is produced due to that bound but for the SnV- as it becomes apparent in Figure 3c). In the Appendix, Sec. 6 we list microwave and laser powers and total processing times for various parameter settings.
In Figure 3c) we illustrate the infidelity $1-\langle F\rangle$ as a function of the upper bound of the cooperativity $C_{\rm max}$ for the optimization of a cavity for the SnV- in a low strain environment at $T=0.1$ K for $\gamma_{\rm in,out}=1$ GHz at the magnetic field strength $B_{\rm dc}=300$ mT and orientation $\theta_{\rm dc}=\pi/2$ using microwave control with $B_{\rm ac}=1$ mT and the orientation $\theta_{\rm ac}=0$. We observe that the infidelity decreases as the cooperativity bound increases. That is the expected behavior since a higher cooperativity yields a flatter shape of the phase contrast in a local environment of the optimized central frequency of the incoming photonic mode [56].
In Figure 3d) we illustrate the infidelity as a function of the infidelity of the auxiliary photon source for read-out assuming the SiV- as the defect center with $B_{\rm dc}=300$ mT, $B_{\rm ac}=3.7$ mT, $\theta_{\rm dc}=\pi/2$, $\theta_{\rm ac}=0$, $E_{x}=4\cdot 10^{-5}$, $\epsilon_{xy}=0$ and $\gamma_{\rm in,out}=0.1$ GHz. We observe a linear rise of the infidelity as the infidelity of the photon source increases. That is intuitive because the quality of the auxiliary photon is essential to produce high fidelity read-out. However, we observe $1-\langle F\rangle\approx 0.2$ at $1-F_{\rm aux}=0.3$ which occurs because the error in the photon gets washed out when performing measurement of the spin.
Finally, we provide an example in which the user supplies experimental cavity data together with the emission frequency of the photonic mode after frequency conversion for the read-in and read-out processes. We consider an SnV- with a DC magnetic field of $B_{\rm dc}=300$ mT oriented orthogonally to its symmetry axis in a low-strain environment, and we assume a photon bandwidth of $\gamma_{\rm in,out}=0.1$ GHz for both operations. Using the software module [49], the user first determines an optimized cavity configuration and central photon frequency specified by $(k,\delta_{c},\delta_{0})$, which then serves as the experimental design target. The software returns the theoretical optimal parameters $\kappa_{\rm theo}=42.85$ GHz, $\delta_{c,\rm theo}=-7.67$ GHz, and $\delta_{0,\rm theo}=-3.41$ GHz. Experimentally, however, the user measures $\kappa_{\rm exp}=41$ GHz, $\delta_{c,\rm exp}=-4$ GHz, and $\delta_{0,\rm exp}=-3$ GHz. We evaluate the average fidelity and trace reduction on the Bloch sphere (see Appendix, Sec. 5) for both the theoretical and experimental parameter sets. The theoretical values are $\langle F\rangle=0.9898$ and $\langle{\rm tr}(\rho_{\rm ph})\rangle=0.9688$, whereas the experimentally inferred values are $\langle F\rangle=0.8306$ and $\langle{\rm tr}(\rho_{\rm ph})\rangle=0.9779$. Thus, although the trace remains nearly unchanged, the fidelity decreases by about $0.15$, indicating that the experimental configuration leaves considerable room for improvement to achieve a high-quality spin memory.
The results indicate that the developed software performs reliably across a wide range of parameters, validating its practical utility.